Grade 12Physical Chemistry

The question says, at infinite dilution (^°) weak mono basic HA axid = 390.7 S cm^2 mol^-1 (^m) of HA at 0.01 M is 3.907 S cm^2 mol^-1 Find the pH of 0.01 M HA?

Shilpi Mazumdar
7 Years agoGrade 12

1 Answer

Sahani Kumar
7 Years ago
Given:
Concentration (ć) =0.01
(At c=0.01)Molar conductivity =3.907
(At infinite dilution)Molar conductivity =390.7
    Now we know that 
degree of dissociation (α)=(molar conductivity at some concentration)/(molar conductivity at infinite dilution) 
So α=(3.907)/(390.7)=10^-2
 
 
Concentration of H+=c*α           [H+]=c*α
                                                       [H+]=0.01*(10^-2) 
                                                       [H+]=10^-4
Now pH= – log[H+]
         pH= – log[10^-4]
         10^-pH = 10^-4
                 pH=4